In Desmos, type in the coordinates:

(1,3) (9,2)

(3,1)

(0,1) (, 1,0)

(-1, 0.333) (0.333, -1)


What is their relationship between x and y coordinates?

1
Expert's answer
2021-05-07T14:39:23-0400

The given points are-

(1,3) (9,2)


The graph is-




Relation between x and y: x+8y=25


The given points are-

(3,1)


The graph is-




Relation between x and y: x=3y


The given points are-

(0,1),(1,0)


The graph is-




Relation between x and y: x + y=1


The given points are-

(-1,0.333),(0.333,-1)


The graph is-





Relation between x and y : x+y+0.66=0